OpenStudy (anonymous):

The heat H in watts generated by an electric heating element varies directly as the square of the current i in amperes. A heating element produces thermal energy at a rate of 648 watts when the current is 0.9 amperes. What rate of heating in watts will be generated by a current of 1.4 amperes?

OpenStudy (anonymous):

Set up a simple equation. We'll call the unknown "r" ;) \[H = i^2r\]

OpenStudy (anonymous):

You know H, you know i. Solve for r, and plug it into the same equation knowing i = 1.4 and r.

OpenStudy (anonymous):

is the answer 800?

OpenStudy (anonymous):

Yes, in this case "r" would be 800. Now plug 800 back into the same equation w/ i = 1.4 \[H = (1.4^2)(800)\] That result will be the solution.

OpenStudy (anonymous):

1568?

OpenStudy (anonymous):

You got it.

OpenStudy (anonymous):

you are so awesome thank you so much. I have more!! :)